# DDR: Exploiting Deep Degradation Response as Flexible Image Descriptor

### v2

@@ 1 Introduction
We evaluate the performance of our proposed DDR across multiple downstream tasks. Firstly, we assess its effectiveness as an image quality descriptor on the blind image quality assessment (BIQA) task, where DDR demonstrates superior performance compared to existing opinion-unaware BIQA methods across various datasets. Secondly, we employ DDR as an unsupervised learning objective, training image restoration models specifically to maximize the DDR of the output image, which includes tasks such as image deblurring and real-world single-image super-resolution. Incorporating DDR as an external training objective consistently improves performance in both tasks, highlighting the strength of DDR as a flexible and powerful image descriptor.
@@ 4.1 Experiment Setting
To demonstrate the versatility and efficacy of DDR as an image descriptor, we conduct comprehensive experiments covering (1) blind image quality assessment (BIQA) and (2) image restoration tasks, including image deblurring and real-world image super-resolution.
Implementation Details. For different tasks, we tailor the degradation set in Eq. 7 and Eq. 8 to focus on distinct image attributes. Specifically, in BIQA, we define . Meanwhile, for image restoration tasks, we set . In all experiments related to image restoration, we empirically set the weight of DDR in the learning objective in Eq. 8 as . We use the CLIP [37] ViT-B/32 model as the image feature extractor. Finally, we employ the cosine distance to quantify the disparity between original and degraded image features.
@@
Baseline Methods. For the BIQA task, we compare DDR with representative and state-of-the-art opinion-unaware BIQA (OU-BIQA) methods, which do not require training with human-labeled MOS. The compared methods include NIQE [26], QAC [55], PIQE [56], LPSI [57], ILNIQE [15], diqIQ [58], SNP-NIQE [59], NPQI [60], and ContentSep [61]. Among all compared methods, DDR is the only zero-shot method that does not require any training. For image restoration, we compare our proposed method, as illustrated in Eq. 8, with a combination of reconstruction loss and feature domain loss , which quantifies the distance between deep features extracted from images. Generally, the reconstruction loss is combined with feature domain losses to enhance the overall quality of the restored image, forming the learning objective of the restoration model as follows:
where is the weighting factor for . In all experiments on image restoration, we utilize PSNR loss as the reconstruction loss. We consider four types of representative feature domain losses for comparison, including LPIPS [3], CTX [2], PDL [7], and FDL [8]. To ensure a fair comparison, we set for FDL [8] and for the other feature domain losses, ensuring that the magnitudes of the different feature domain losses are in a similar range. To fully assess the robustness of our proposed DDR across various architectural models, we conduct all image restoration experiments using two representative image restoration models: NAFNet [62] and Restormer [63]. NAFNet [62] is a convolutional neural network (CNN)-based model, while Restormer [63] is a Transformer [64]-based model. These models have demonstrated impressive performance in their respective tasks and are widely recognized as representative models in recent years.
4.2 Blind Image Quality Assessment
@@ 4.5 Ablation Study
For image deblurring, we conduct a series of ablation experiments on the NAFNet using the GoPro dataset, with all results detailed in Table 6. Additionally, further results from ablation experiments for blind image quality assessment are provided in the appendix. Firstly, we adjusted to evaluate the effect of the degradation set defined in Eq. 8. Notably, a decrease in performance is observed when any type of degradation is removed, suggesting that combining multiple types of degradation results in a more comprehensive image description.
Secondly, we investigate the effect of in Eq. 8. Minor fluctuations in performance are observed when adjusting to 1.0 and 3.0, indicating the robustness of our method to this hyper-parameter. Next, we explore the effect of the visual feature extractor. Increasing the scale of in DDR does not lead to improved performance, suggesting that a larger visual model may not necessarily enhance the ability to understand low-level texture. Finally, we examine the impact of the adaptation strategy in Eq. 4. A significant drop in performance is observed when the adaptation is removed, underscoring the importance of this strategy in the calculation of DDR.

### v3

@@ 1 Introduction
We evaluate the performance of our proposed DDR across multiple downstream tasks. Firstly, we assess its effectiveness as an image quality descriptor on the opinion-unaware blind image quality assessment (OU-BIQA) task, where DDR demonstrates superior performance compared to existing OU-BIQA methods across various datasets. Secondly, we employ DDR as an unsupervised learning objective, training image restoration models specifically to maximize the DDR of the output image, which includes tasks such as image deblurring and real-world single-image super-resolution. Incorporating DDR as an external training objective consistently improves performance in both tasks, highlighting the strength of DDR as a flexible and powerful image descriptor.
@@ 4.1 Experiment Setting
To demonstrate the versatility and efficacy of DDR as an image descriptor, we conduct comprehensive experiments covering (1) opinion-unaware blind image quality assessment (OU-BIQA), which does not require training model with human-labeled Mean Opinion Score (MOS) values, and (2) image restoration tasks, including image deblurring and real-world image super-resolution.
Implementation Details. For different tasks, we tailor the degradation set in Eq. 7 and Eq. 8 to focus on distinct image attributes. Specifically, in BIQA, we define . Meanwhile, for image restoration tasks, we set . In all experiments related to image restoration, we empirically set the weight of DDR in the learning objective in Eq. 8 as . We use the CLIP [37] ViT-B/32 model as the image feature extractor, and employ the cosine distance to quantify the disparity between original and degraded image features, which can be defined as follows:
@@
Baseline Methods. For the OU-BIQA task, we compare DDR with representative and state-of-the-art opinion-unaware BIQA (OU-BIQA) methods, which do not require training with human-labeled MOS. The compared methods include NIQE [26], QAC [55], PIQE [56], LPSI [57], ILNIQE [15], diqIQ [58], SNP-NIQE [59], NPQI [60], and ContentSep [61]. Among all compared methods, DDR is the only zero-shot method that does not require any training. For image restoration, we compare our proposed method, as illustrated in Eq. 8, with a combination of reconstruction loss and feature domain loss , which quantifies the distance between deep features extracted from images. Generally, the reconstruction loss is combined with feature domain losses to enhance the overall quality of the restored image, forming the learning objective of the restoration model as follows:
where is the weighting factor for . In all experiments on image restoration, we utilize PSNR loss as the reconstruction loss. We consider four types of representative feature domain losses for comparison, including LPIPS [3], CTX [2], PDL [7], and FDL [8]. To ensure a fair comparison, we set for FDL [8] and for the other feature domain losses, ensuring that the magnitudes of the different feature domain losses are in a similar range.
Moreover, to fully assess the robustness of our proposed DDR across vaious architectural models, we conduct all image restoration experiments using two representative image restoration models: NAFNet [62] and Restormer [63]. NAFNet [62] is a convolutional neural network (CNN)-based model, while Restormer [63] is a Transformer [64]-based model. These models have demonstrated impressive performance in their respective tasks and are widely recognized as representative models in recent years. We empirically train the model at a resolution of . For the learning rate, we adhere to the official settings for NAFNet and Restormer. Specifically, the initial learning rate for NAFNet is set to , and for Restormer, it is set to . We also adopted a cosine annealing strategy for both models.
4.2 Opinion-Unaware Blind Image Quality Assessment
@@ 4.5 Ablation Study
For image deblurring, we conduct a series of ablation experiments on the NAFNet using the GoPro dataset, with all results detailed in Table 6. Firstly, we adjusted to evaluate the effect of the degradation set defined in Eq. 8. Notably, a decrease in performance is observed when any type of degradation is removed, suggesting that combining multiple types of degradation results in a more comprehensive image description. Secondly, we investigate the effect of in Eq. 8. Minor fluctuations in performance are observed when adjusting to 1.0 and 3.0, indicating the robustness of our method to this hyper-parameter. Next, we explore the effect of the visual feature extractor. Increasing the scale of in DDR does not lead to improved performance, suggesting that a larger visual model may not necessarily enhance the ability to understand low-level texture. Finally, we examine the impact of the adaptation strategy in Eq. 4. A significant drop in performance is observed when the adaptation is removed, highlighting the critical role of this strategy in DDR calculation.
For opinion-unaware blind image quality assessment task, we conduct ablation experiment on four datasets. As demonstrated in Tab. 9, comparing with measuring DDR to single type of degradation, combining mutiple degradation types consistently leads to significant performance improvement. Furthermore, we can observe a performance boost by utilizing degradation adaptation strategy, improving SRCC from 0.6074 to 0.8289 on CSIQ dataset.
| Backbone | Adaptation | CSIQ | TID2013 | LIVEitw | KonIQ | |
| ViT-B/32 | ✓ | 0.7896 | 0.4872 | 0.5178 | 0.5745 | |
| ViT-B/32 | ✓ | 0.7555 | 0.5061 | 0.6282 | 0.6028 | |
| ViT-B/32 | ✓ | 0.6937 | 0.5022 | 0.4954 | 0.5804 | |
| ViT-B/32 | ✓ | 0.7695 | 0.5440 | 0.6312 | 0.5758 | |
| ViT-B/32 | ✓ | 0.8029 | 0.5841 | 0.6660 | 0.6390 | |
| ViT-B/32 | ✓ | 0.8100 | 0.5917 | 0.6352 | 0.6379 | |
| ViT-B/32 | ✓ | 0.8054 | 0.5446 | 0.6548 | 0.6229 | |
| ViT-B/32 | ✓ | 0.8157 | 0.5824 | 0.6399 | 0.6526 | |
| ViT-B/32 | ✗ | 0.6074 | 0.4121 | 0.3437 | 0.4310 | |
| ViT-B/16 | ✓ | 0.7841 | 0.6251 | 0.6888 | 0.6635 | |
| ViT-B/32 | ✓ | 0.8289 | 0.5844 | 0.6613 | 0.6455 | |
| RN50x16 | ✓ | 0.6500 | 0.5464 | 0.6607 | 0.6125 | |
| RN50x64 | ✓ | 0.6162 | 0.5607 | 0.6703 | 0.6109 |
5 Limitations and Discussion
In this section, we discuss the limitations of DDR and provide potential solutions to address them.
The ability of the visual feature extractor to understand low-level degradation. We currently employ the CLIP model’s visual feature extractor to facilitate text-driven degradation fusion. These feature extractors may incline to focus on high-level information such as image content, while their ability to understand low-level degradation may be limited. This could impact the measurement of the degradation response. In future work, we plan to fine-tune the feature extractor on tasks such as degradation classification or description, to enable it to extract more fine-grained degradation features.